OpenStudy (anonymous):

simplify [ (3x^(3/2)Y^3)/(x^2y^-1/2) ]^2

OpenStudy (gorv):

\[[\frac{ 3*x^{3/2} *y^3}{ (x^{2}*y^{-1/2}) }]^{2}\]

OpenStudy (anonymous):

sorry its all to the -2

OpenStudy (gorv):

\[[3*x^{\frac{ 3 }{ 2 }-2}*y^{3-\frac{ (-1) }{ 2 }}]^{-2}\]

OpenStudy (anonymous):

\[3x \frac{3y^{3} }{2 } {-4} y\]

OpenStudy (gorv):

\[[3*x^{\frac{ -1 }{ 2 }}*y^{\frac{ 7 }{ 2 }}]^{-2}\]

OpenStudy (gorv):

\[3^{-2}*x^{(\frac{ -1 }{ 2 }*(-2))}*y^{(\frac{ 7 }{ 2 }*(-2))}=\frac{ 1 }{ 9 }*x^{1}*y^{-7}\]

OpenStudy (gorv):

\[\frac{ x }{ 9*y ^{7} }\] @Ssusu

OpenStudy (anonymous):

thank uu @gorv

OpenStudy (anonymous):

\[3x \frac{ 11 }{ 2 }y2\]

OpenStudy (gorv):

@Ssusu always welcme :)